Vinny purchased an annuity contract for $153,300 on April 30, 20X1. He is to receive $700 per month starting June 1, 20X1, and continuing for life. He has a life expectancy of 25 years. What is the amount of Vinny's annuity income that must be included in his gross income for 20X1?

a. $3,577

b. $1,323

C. $6,132

d. $2,268

e. None of the above
Answer #1
Total amount to be received 210000 =700*25*12
Less: Cost of Annuity 153300
Taxable annuity 56700
Taxable annuity per month 189 =56700/(25*12)
Vinny will receive 7 payments in 20X1 (June-December)
Amount included in Gross income for 20X1 1323 =189*7
Option B $1,323 is correct